OpenStudy (anonymous):

every +ve odd integer is of the form: (a) 2m (b) m (c) 2m-1 (d) 2m+1

OpenStudy (anonymous):

Both c and d are possible. \( 2m-1 ,m\ge 1, m\in \mathbb{N} \) \( 2m+1, m\ge 0, m\in \mathbb{W} \)

OpenStudy (anonymous):

wch is d option